OpenStudy (anonymous):

Please Help! Show your work including the tables and graphs. 1. Using complete sentences, explain how to find the zeros of the function f(x) = 2x^3 – 9x + 3

OpenStudy (cwrw238):

you can first draw a graph to find the intercepts on x-axis there will be up to 3 roots in this cubic polynomial one will be close to 2 because 2(2)^3 - 9(2) + 3 = 16 - 18 + 3 = 1 use a graphical computer or the wolframalpha site or you could plot points i suggest you plot points from x = -5 to x = 5

OpenStudy (anonymous):

i still dont get i. i graphed the equation and this is what i got..

OpenStudy (cwrw238):

looks good you have 3 zeros where the graph crosses the x-axis